OpenStudy (anonymous):

how do i find the range and the min and max values? 13−20x−x^2−3x^4

OpenStudy (amistre64):

calculus helps

OpenStudy (amistre64):

from algebra, we know that a -x^4 graph does what?

OpenStudy (anonymous):

goes through the x-axis 4 times?

OpenStudy (amistre64):

at least 4 times, but in terms of range, it opens downward, so the max will approach -inf

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

the min, we should take a derivative

OpenStudy (anonymous):

i havent learned how to do that yet. this is from a summer packet that i have to do in preparation for AP calculus

OpenStudy (amistre64):

13−20x−x^2−3x^4 derive to get -20-2x-12x^3 and determine when -20-2x-12x^3=0

OpenStudy (amistre64):

your packet doesnt tell you how to find a derivative? or a min/max of a function?

OpenStudy (anonymous):

No, i havent learned derivatives yet, is there any other way i could approach this?

OpenStudy (amistre64):

trial and error ..

OpenStudy (anonymous):

what if i just graphed it?

OpenStudy (amistre64):

if you are able to just graph it, then yeah that would be fine

OpenStudy (anonymous):

okay thank you

OpenStudy (amistre64):

-20-2x-12x^3=0 -2(10+x+6x^3) = 0 10+x+6x^3 = 0 has no positive roots and 1 negative root, which would have to be a double so im thinking the graph is something like: